# PlanGenLLMs: A Modern Survey of LLM Planning Capabilities

### v1

@@ Abstract
LLMs have immense potential for generating plans, transforming an initial world state into a desired goal state. A large body of research has explored the use of LLMs for various planning tasks, from web navigation to travel planning and database querying. However, many of these systems are tailored to specific problems, making it challenging to compare them or determine the best approach for new tasks. There is also a lack of clear and consistent evaluation criteria. Our survey aims to offer a comprehensive overview of current LLM planners to fill this gap. It builds on foundational work by Kartam and Wilkins (1990) and examines six key performance criteria: completeness, executability, optimality, representation, generalization, and efficiency. For each, we provide a thorough analysis of representative works and highlight their strengths and weaknesses. Our paper also identifies crucial future directions, making it a valuable resource for both practitioners and newcomers interested in leveraging LLM planning to support agentic workflows.
@@ 1 Introduction
LAST Zhou et al. (2023a); Valmeekam et al. (2023b); SimPlan Hirsch et al. (2024); Tree Search Koh et al. (2024). ,leaf ,text width=55.5em] ] [Search Algorithm,text width=10.9em [ Tree of Thought Yao et al. (2024); Thought of Search Katz et al. (2024); LLM-MCTS Zhao et al. (2024);
RAP Hao et al. (2023); LAST Zhou et al. (2023a); SimPlan Hirsch et al. (2024); Tree Search Koh et al. (2024);
@@
     (§3 - §8) [Completeness (§3) [Plan Correctness ,text width=13.4em [Guan et al. (2023); Hao et al. (2024a). , leaf, text width= 42.5em] ] [Plan Achievability,text width=13.4em [PPNL Aghzal et al. (2023); Valmeekam et al. (2024). ,leaf, text width = 42.5em] ] ] [Executability (§4) [Object Grounding,text width=13.4em [AdaPlanner Sun et al. (2024); Inner Monologue Huang et al. (2022b);
@@
G-PlanE Lin et al. (2023a). ,leaf, text width=42.5em] ] [Action Grounding,text width=13.4em [LLM-Planner Song et al. (2023); BrainBoday-LLM Bhat et al. (2024);
Corrective Re-prompting Raman et al. (2022); SayCanPay Hazra et al. (2024). ,leaf, text width= 42.5em] ] [Sample-then-Filter,text width=13.4em [CLP Yuan et al. (2023); PLASMA Brahman et al. (2024); PRoC3S Curtis et al. (2024). ,leaf, text width= 42.5em] ] [Closed-Loop Systems,text width=13.4em [Corrective Re-prompting Raman et al. (2022); ProgPrompt Singh et al. (2023);
@@
ISR-LLM Zhou et al. (2024). ,leaf, text width= 42.5em] ] ] [Optimality (§5) [LLM + Optimizer,text width=13.4em [TTG Ju et al. (2024); LLMFP Hao et al. (2024b). ,leaf, text width= 42.5em] ] [A* Search-Based Methods,text width=13.4em [ToolChain* Zhuang et al. (2023); SayCanPay Hazra et al. (2024);
Beyond A* Lehnert et al. (2024). ,leaf, text width= 42.5em] ] ] [Representation (§6) [LLM-as-a-Translator,text width=13.4em [LLM-GoalTrans Xie et al. (2023); ISR-LLM Zhou et al. (2024); Adaplanner Sun et al. (2024);
@@
LLM-DP Dagan et al. (2023); Guan et al. (2023). ,leaf, text width= 42.5em] ] [LLM-as-a-Planner,text width=13.4em [G-PlanET Lin et al. (2023a); Chain-of-symbols Hu et al. (2024); PPNL Aghzal et al. (2023);
@@
ProgPrompt Singh et al. (2023); LID Li et al. (2022). ,leaf, text width= 42.5em] ] ] [Generalization (§7) [Fine-tuning,text width=13.4em [ RobLM Chalvatzaki et al. (2023); Agent-FLAN Chen et al. (2024b); ETO Song et al. (2024);
AgentOhana Zhang et al. (2024); NAT Wang et al. (2024); Jansen (2020). ,leaf ,text width= 42.5em] ] [Generalized Planning,text width=13.4em [ LLM-GenPlan Silver et al. (2024). ,leaf ,text width= 42.5em] ] [Skill Storage,text width=13.4em [VOYAGER Wang et al. (2023a). ,leaf ,text width= 42.5em] ] ] [Efficiency (§8) [Reduced LLM and
World Model Calls,text width=13.4em [AdaPlanner Sun et al. (2024); Query-Efficient Planning Gonzalez-Pumariega et al. (2024);
Thought of Search Katz et al. (2024); Tree-Planner Hu et al. (2023b). ,leaf ,text width= 42.5em] ] [Shorter Inputs and Outputs,text width=13.4em [Chain-of-Symbols Hu et al. (2024); Beyond A* Lehnert et al. (2024);
DEPS Wang et al. (2023b); LLM-DP Dagan et al. (2023). ,leaf ,text width= 42.5em] ] [Smaller Model Size,text width=13.4em [PLASMA Brahman et al. (2024). ,leaf ,text width= 42.5em] ] ] ] [  Evaluation
@@
Tool Usage: ToolBench Xu et al. (2023); APIBench Patil et al. (2023); TPTU Ruan et al. (2023).
@@ 2 LLM Planning Foundations (Tables 1-3)
Current LLMs are not specifically trained for agentic tasks like planning, and prompt-based methods, which do not update model parameters, cannot fundamentally improve performance in these areas Chen et al. (2023a); Wang et al. (2024). Fine-tuning, either focused on planning-specific tasks or broader agentic capabilities, enhances planning correctness by directly updating LLM parameters.
@@ 3 Criterion I: Completeness (Table 4)
A plan is correct if it achieves the goal within a fixed budget while avoiding excessive complexity and infinite loops. To ensure correctness, the LLM must work with classical sound and complete solvers Guan et al. (2023); Hao et al. (2024a). Also, the LLM has to accurately translate the domain and problem into the specific format (e.g., PDDL), required by these solvers Guan et al. (2023).
In terms of identifying unsolvable planning problems, those with inherently unachievable goals, even top LLMs (e.g., GPT-4 Achiam et al. (2023)) and Large Reasoning Models (e.g., OpenAI O1 Jaech et al. (2024)) struggle due to hallucination issues Aghzal et al. (2023); Valmeekam et al. (2024).
@@ 4 Criterion II: Executability (Tables 5-6)
Object Grounding Object grounding ensures the LLM planner uses objects available in the current environment when generating plans. The simplest way to do this is by feeding observed or available objects into the planner via prompts (Huang et al., 2022b; Song et al., 2023; Lin et al., 2023a; Singh et al., 2023) or neural embeddings Sharma et al. (2021); Ahn et al. (2022). In partially observed environments, where some object information are uncertain (e.g., needing to clean a cup that could be in a cabinet, drawer, or fridge), the planner can generate multiple possible plans, one for each scenario, and select the first feasible one (Prasad et al., 2023; Dagan et al., 2023; Zhao et al., 2024). Sun et al. (2024) takes a different approach, first generating a plan with placeholders for objects, then filling in the blanks with observed objects during execution.
Action Grounding Action grounding ensures all actions in a plan can actually be executed in the current environment. Like object grounding, the simplest way is to explicitly list all admissible actions in LLMs’ inputs Singh et al. (2023). If a step goes beyond the executor’s capabilities (e.g., combining multiple allowed actions into one), the LLM planner should be reprompted to break it down until every step is executable (Prasad et al., 2023).
@@
Closed-Loop Systems A closed-loop system in LLM planning means the model adapts its plan based on feedback from executors Prasad et al. (2023); Yang et al. (2024), simulators Bhat et al. (2024), validators Zhou et al. (2024); Silver et al. (2024), other LLMs Wang et al. (2023b), or even humans Huang et al. (2022b), when the initial plan are inexecutable. It reprompts the LLM planner to replan until the plan is fully executable. Unlike open-loop systems Huang et al. (2022a), which lack feedback, closed-loop planning helps reduce hallucinations and enables LLMs to handle complex, long-horizon, and dynamic environments Wang et al. (2023b).
@@ 5 Criterion III: Optimality (Table 7)
LLM + Optimizer   It combines the LLM, which turns user requests into symbolic optimization problems, with an optimizer that solves them and finds the best solution Ju et al. (2024); Hao et al. (2024b). For example, TTG (Ju et al., 2024) uses the LLM to convert travel planning requests of minimum total costs into Mixed Integer Linear Programming problems, then runs an optimizer such as SCIP (Bestuzheva et al., 2021) to provide the optimal plan. Compared to LLM + classical planners, where LLMs define the domain and problem in a formal representation, LLM + optimizers ensure optimal solutions by further formulating and solving constrained optimization problems.
A* Search-Based Methods A* search always finds the lowest-cost optimal solution, making it a natural choice for LLM-based planners to achieve optimality. ToolChain* Zhuang et al. (2023) combines A* tree search with an LLM, which suggests next steps and estimates heuristic scores, to create plans with the fewest tool API calls. SayCanPay (Hazra et al., 2024) uses A* search with LLMs to generate the shortest possible plans. Beyond A* (Lehnert et al., 2024) trains a Transformer model, Searchformer, to mimic A* search paths for complex tasks like Maze navigation and Sokoban puzzles, optimizing for the fewest steps. Besides A* search, other search algorithms (e.g., DFS and MCTS) can also be used to find optimal solutions.
@@
Recent work has used LLMs to translate natural language descriptions into PDDL Liu et al. (2023a); Guan et al. (2023); Xie et al. (2023); Dagan et al. (2023); Zhou et al. (2024), LTL Pan et al. (2023), and STL Chen et al. (2024a). To ensure reliability, translations should be tested on development or external datasets like Planetarium Zuo et al. (2024). If there are syntax or semantic errors, validators (e.g. VAL Howey et al. (2004)) or human experts can provide feedback for the LLM to fix them.
@@ 7 Criterion V: Generalization (Table 10)
Generalized planning extracts common patterns from training solutions to tackle larger, more complex tasks within the same domain Srivastava et al. (2011). For example, in the Delivery dataset Yang et al. (2022), models trained on small-scale deliveries (9–17 locations) can generalize to larger ones (70–100 locations) using the same core strategy. Silver et al. (2024) approached this by prompting LLMs to summarize the domain and generate a minimal, generalizable Python-based plan.
@@ 8 Criterion VI: Efficiency (Table 11)
To decrease LLM and world model calls, several tricks are used: (1) generating the entire plan in one shot instead of step-by-step to reduce redundant prompts Hu et al. (2023b); Sun et al. (2024); Gonzalez-Pumariega et al. (2024); (2) checking feasibility by world models only at the end of each subgoal, not after every action (Sun et al., 2024; Gonzalez-Pumariega et al., 2024); (3) merging plans with the same prefix actions or subgoals to avoid duplicate world model checks when sampling multiple plans (Hu et al., 2023b); and (4) in tree search-based methods, querying the LLM once to list all possible actions and states, and generate a state transition and goal check function, avoiding repeated LLM and world model calls at each node (Katz et al., 2024).
@@ 9 Evaluation
Planning-focused datasets primarily assess planning abilities. The most common scenarios include (1) Embodied environments, (2) Task scheduling , (3) Games, and (4) Task decomposition Li et al. (2024d). Figure 1 presents commonly used planning datasets; readers can refer to Li et al. (2024d) for further details.
@@
Planning abilities can also be evaluated through downstream tasks, where planning is integral to task completion, and stronger planning skills enhance overall performance. Downstream tasks can be categorized as follows: (1) Agentic tasks, including reasoning-oriented tasks, tool-use-oriented tasks, programming tasks, and web tasks Zhou et al. (2023b); Liu et al. (2023b); Li et al. (2023); Xu et al. (2023), (2) Generation tasks, including video Lin et al. (2023b), image Zala et al. (2023) and text generation Moryossef et al. (2019). Please refer to Figure 1 for example datasets.

### v2

@@ Abstract
LLMs have immense potential for generating plans, transforming an initial world state into a desired goal state. A large body of research has explored the use of LLMs for various planning tasks, from web navigation to travel planning and database querying. However, many of these systems are tailored to specific problems, making it challenging to compare them or determine the best approach for new tasks. There is also a lack of clear and consistent evaluation criteria. Our survey aims to offer a comprehensive overview of current LLM planners to fill this gap. It builds on foundational work by Kartam and Wilkins (1990) and examines six key performance criteria: completeness, executability, optimality, representation, generalization, and efficiency. For each, we provide a thorough analysis of representative works and highlight their strengths and weaknesses. Our paper also identifies crucial future directions, making it a valuable resource for both practitioners and newcomers interested in leveraging LLM planning to support agentic workflows. [^1] A curated list of papers and resources related to this survey are available at https://github.com/wll199566/Awesome-LLM-Planning-Capability.
@@
LATS Zhou et al. (2023a); Valmeekam et al. (2023b); SimPlan Hirsch et al. (2024); Tree Search Koh et al. (2024). ,leaf ,text width=55.5em] ] [Search Algorithm,text width=10.9em [ Tree of Thought Yao et al. (2024); Thought of Search Katz et al. (2024); LLM-MCTS Zhao et al. (2024);
RAP Hao et al. (2023); LATS Zhou et al. (2023a); SimPlan Hirsch et al. (2024); Tree Search Koh et al. (2024);
@@
     (§3 - §8) [Completeness (§3) [Plan Correctness ,text width=12.4em [Guan et al. (2023); Hao et al. (2024a). , leaf, text width= 43.5em] ] [Plan Achievability,text width=12.4em [PPNL Aghzal et al. (2023); Valmeekam et al. (2024). ,leaf, text width = 43.5em] ] ] [Executability (§4) [Object Grounding,text width=12.4em [AdaPlanner Sun et al. (2024); Inner Monologue Huang et al. (2022b);
@@
G-PlanE Lin et al. (2023a). ,leaf, text width=43.5em] ] [Action Grounding,text width=12.4em [LLM-Planner Song et al. (2023); BrainBoday-LLM Bhat et al. (2024);
Corrective Re-prompting Raman et al. (2022); SayCanPay Hazra et al. (2024). ,leaf, text width= 43.5em] ] [Sample-then-Filter,text width=12.4em [CLP Yuan et al. (2023); PLASMA Brahman et al. (2024); PRoC3S Curtis et al. (2024). ,leaf, text width= 43.5em] ] [Closed-Loop Systems,text width=12.4em [Corrective Re-prompting Raman et al. (2022); ProgPrompt Singh et al. (2023);
@@
ISR-LLM Zhou et al. (2024). ,leaf, text width= 43.5em] ] ] [Optimality (§5) [LLM + Optimizer,text width=12.4em [TTG Ju et al. (2024); LLMFP Hao et al. (2024b). ,leaf, text width= 43.5em] ] [A* Search-Based Methods,text width=12.4em [ToolChain* Zhuang et al. (2023); SayCanPay Hazra et al. (2024);
Beyond A* Lehnert et al. (2024). ,leaf, text width= 43.5em] ] ] [Representation (§6) [LLM-as-a-Translator,text width=12.4em [LLM-GoalTrans Xie et al. (2023); ISR-LLM Zhou et al. (2024); Adaplanner Sun et al. (2024);
@@
LLM-DP Dagan et al. (2023); Guan et al. (2023). ,leaf, text width= 43.5em] ] [LLM-as-a-Planner,text width=12.4em [G-PlanET Lin et al. (2023a); Chain-of-symbols Hu et al. (2024); PPNL Aghzal et al. (2023);
@@
ProgPrompt Singh et al. (2023); LID Li et al. (2022). ,leaf, text width= 43.5em] ] ] [Generalization (§7) [Fine-tuning,text width=12.4em [ RobLM Chalvatzaki et al. (2023); Agent-FLAN Chen et al. (2024b); ETO Song et al. (2024);
AgentOhana Zhang et al. (2024); NAT Wang et al. (2024); Jansen (2020). ,leaf ,text width= 43.5em] ] [Generalized Planning,text width=12.4em [ LLM-GenPlan Silver et al. (2024). ,leaf ,text width= 43.5em] ] [Skill Storage,text width=12.4em [VOYAGER Wang et al. (2023a). ,leaf ,text width= 43.5em] ] ] [Efficiency (§8) [Reduced LLM and
World Model Calls,text width=12.4em [AdaPlanner Sun et al. (2024); Query-Efficient Planning Gonzalez-Pumariega et al. (2024);
Thought of Search Katz et al. (2024); Tree-Planner Hu et al. (2023b). ,leaf ,text width= 43.5em] ] [Shorter Inputs and Outputs,text width=12.4em [Chain-of-Symbols Hu et al. (2024); Beyond A* Lehnert et al. (2024);
DEPS Wang et al. (2023b); LLM-DP Dagan et al. (2023). ,leaf ,text width= 43.5em] ] [Smaller Model Size,text width=12.4em [PLASMA Brahman et al. (2024). ,leaf ,text width= 43.5em] ] ] ] [  Evaluation
@@
Tool Usage: ToolBench Xu et al. (2023); API-Bank Li et al. (2023); TPTU Ruan et al. (2023).
@@ 2 LLM Planning Foundations (Tables 1-3)
Commonly used pretrained LLMs (e.g., GPT-4 Achiam et al. (2023), Claude 3 Anthropic (2024), Llama 2 Touvron et al. (2023)) are not specifically trained for planning tasks, and prompt-based methods, which do not update model parameters, cannot fundamentally improve performance in these areas Chen et al. (2023a); Wang et al. (2024). Fine-tuning, either focused on planning-specific tasks or broader agentic capabilities, enhances planning correctness by directly updating LLM parameters.
@@ 3 Criterion I: Completeness (Table 4)
A plan is correct if it achieves the goal within a fixed budget while avoiding excessive complexity and infinite loops (e.g., finishing booking a flight ticket to San Francisco within five hours meets the criteria for correctness). To ensure correctness, the LLM must work with classical sound and complete solvers Guan et al. (2023); Hao et al. (2024a). Also, the LLM has to accurately translate the domain and problem into the specific format (e.g., PDDL), required by these solvers Guan et al. (2023).
In terms of identifying unsolvable planning problems, those with inherently unachievable goals, even top LLMs (e.g., GPT-4 Achiam et al. (2023)) and Large Reasoning Models (e.g., OpenAI O1 Jaech et al. (2024)) struggle due to hallucination issues Aghzal et al. (2023); Valmeekam et al. (2024); Katz et al. (2024).
@@ 4 Criterion II: Executability (Tables 5-6)
Object Grounding Object grounding ensures the LLM planner uses objects available in the current environment when generating plans. E.g., if a stove is present but a microwave is not, the planner should correctly select the stove to heat a pancake and exclude the unavailable microwave from its plan. The simplest way to do this is by feeding observed or available objects into the planner via prompts (Huang et al., 2022b; Song et al., 2023; Lin et al., 2023a; Singh et al., 2023) or neural embeddings Sharma et al. (2021); Ahn et al. (2022). In partially observed environments, where some object information are uncertain (e.g., needing to clean a cup that could be in a cabinet, drawer, or fridge), the planner can generate multiple possible plans, one for each scenario, and select the first feasible one (Prasad et al., 2023; Dagan et al., 2023; Zhao et al., 2024). Sun et al. (2024) takes a different approach, first generating a plan with placeholders for objects, then filling in the blanks with observed objects during execution.
Action Grounding Action grounding ensures all actions in a plan can actually be executed in the current environment. E.g., the planner should not include an action like “draw a star” if the robot arm cannot perform such a complex operation. Instead, the task should be decomposed into simpler supported actions, such as multiple steps of “draw a line”. Like object grounding, the simplest way is to explicitly list all admissible actions in LLMs’ inputs Singh et al. (2023). If a step goes beyond the executor’s capabilities (e.g., combining multiple allowed actions into one), the LLM planner should be reprompted to break it down until every step is executable (Prasad et al., 2023).
@@
Closed-Loop Systems A closed-loop system in LLM planning means the model adapts its plan based on feedback from executors Prasad et al. (2023); Yang et al. (2024), simulators Bhat et al. (2024), validators Zhou et al. (2024); Silver et al. (2024), other LLMs Wang et al. (2023b); Zhou et al. (2023a), or even humans Huang et al. (2022b), when the initial plan are inexecutable. It reprompts the LLM planner to replan until the plan is fully executable. Unlike open-loop systems Huang et al. (2022a), which lack feedback, closed-loop planning helps reduce hallucinations and enables LLMs to handle complex, long-horizon, and dynamic environments Wang et al. (2023b).
@@ 5 Criterion III: Optimality (Table 7)
LLM + Optimizer   It combines the LLM, which turns user requests into symbolic optimization problems, with an optimizer that solves them and finds the best solution Ju et al. (2024); Hao et al. (2024b). For example, TTG (Ju et al., 2024) uses the LLM to convert travel planning requests of minimum total costs into Mixed Integer Linear Programming problems, then runs an optimizer such as SCIP (Bestuzheva et al., 2021) to provide the optimal plan. Compared to LLM + classical planners, where LLMs define the domain and problem in a formal representation (e.g., PDDL), LLM + optimizers ensure optimal solutions by leveraging LLMs to formulate constrained optimization problems and classical optimizers to solve them. Unlike classical planners, which typically rely on search algorithms, heuristics and logical deductions and may not guarantee optimality Russell and Norvig (2016), optimizers, often using gradient-based methods (e.g., Newton’s methods), can guarantee optimal solutions Boyd and Vandenberghe (2004).
A* Search-Based Methods A* search finds the lowest-cost optimal solution, particularly when using admissible heuristics that do not overestimate the actual cost to the goal. This makes it a natural choice for LLM-based planners to achieve optimality. ToolChain* Zhuang et al. (2023) combines A* tree search with an LLM, which suggests next steps and estimates heuristic scores, to create plans with the fewest tool API calls. SayCanPay (Hazra et al., 2024) uses A* search with LLMs to generate the shortest possible plans. Beyond A* (Lehnert et al., 2024) trains a Transformer model, Searchformer, to mimic A* search paths for complex tasks like Maze navigation and Sokoban puzzles, optimizing for the fewest steps. Besides A* search, other search algorithms (e.g., DFS and MCTS) could also be used to find optimal solutions, although without guarantee.
@@
Recent work has used LLMs to translate natural language descriptions into PDDL Liu et al. (2023a); Guan et al. (2023); Xie et al. (2023); Dagan et al. (2023); Zhou et al. (2024); Tantakoun et al. (2025), LTL Pan et al. (2023), and STL Chen et al. (2024a). To ensure reliability, translations should be tested on development or external datasets like Planetarium Zuo et al. (2024). If there are syntax or semantic errors, validators (e.g. VAL Howey et al. (2004)) or human experts can provide feedback for the LLM to fix them.
@@ 7 Criterion V: Generalization (Table 10)
Generalized planning extracts common patterns from a limited set of training solutions (i.e., plans) to solve unseen tasks within the same domain, which may be larger and more complex than the training tasks Srivastava et al. (2011). For example, in the Delivery dataset Yang et al. (2022), models trained on small-scale deliveries (9–17 locations) can generalize to larger ones (70–100 locations) using the same core strategy. Silver et al. (2024) approached this by prompting LLMs to summarize the domain and generate a minimal, generalizable Python-based plan.
@@ 8 Criterion VI: Efficiency (Table 11)
To decrease LLM and world model calls, several tricks are used: (1) generating the entire plan in one shot instead of step-by-step to reduce redundant prompts Hu et al. (2023b); Sun et al. (2024); Gonzalez-Pumariega et al. (2024); (2) checking feasibility by world models only at the end of each subgoal, not after every action (Sun et al., 2024; Gonzalez-Pumariega et al., 2024); (3) merging plans with the same prefix actions or subgoals to avoid duplicate world model checks when sampling multiple plans (Hu et al., 2023b); and (4) in tree search-based methods, querying the LLM once to generate a successor function and a goal check function. The successor function, which can produce all possible actions and states, provides state transitions based on the current state and selected action. The goal-check function determines whether the current state is the desired final goal. This approach avoids repeated LLM and world model calls at each node (Katz et al., 2024).
@@ 9 Evaluation
Planning-focused datasets primarily assess planning abilities. The most common scenarios include (1) Embodied environments, (2) Task scheduling , (3) Games, and (4) Task decomposition Li et al. (2024d). Figure 1 presents commonly used planning datasets; readers can refer to Li et al. (2024d, 2025) for further details.
@@
Planning abilities can also be evaluated through downstream tasks, where planning is integral to task completion, and stronger planning skills enhance overall performance. Downstream tasks can be categorized as follows: (1) Agentic tasks, including reasoning-oriented tasks, tool-use-oriented tasks, programming tasks, and web tasks Yu et al. (2018); Cobbe et al. (2021); Saparov and He (2022); Zhou et al. (2023b); Deng et al. (2023); Liu et al. (2023b); Li et al. (2023); Xu et al. (2023); Jimenez et al. (2023); Ruan et al. (2023); Bairi et al. (2023); Li et al. (2024e), (2) Generation tasks, including video Lin et al. (2023b), image Zala et al. (2023) and text generation Moryossef et al. (2019). Please refer to Figure 1 for example datasets.
@@ 12 Limitations
Acknowledgements
We thank the reviewers for their insightful feedback, which has helped improve our paper. H.W. and S.P. are supported by a UC Merced Spring 2023 Climate Action Seed Competition Grant, CAHSI-Google Institutional Research Program Award, and F3 R&D GSR Award funded by the US Department of Commerce, Economic Development Administration Build Back Better Regional Challenge. Z.Z. and F.L. are partially supported by NSF CAREER Award #2303655.